# Mangosteen: An Open Thai Corpus for Language Model Pretraining

### v1

@@ 1 Introduction
For language-specific datasets, many data collection projects utilize existing data cleaning pipelines by customizing them to make these pipelines more suitable for their target languages. This ranges from making a small modification in one process, for instance, the Latxa project Etxaniz et al. (2024) for Basque, which adapted the Dolma and Corpus Cleaner v2 pipelines with a specific change in the filtering step, to customizing the whole pipeline. Examples of more extensive customization include the FinGPT Luukkonen et al. (2023) and IndicLLMSuite Khan et al. (2024) projects that built their own dedicated pipelines for Finnish and Indian languages, respectively. For the Thai language, the approach ranges from using a standard language-agnostic pipeline, such as in the Typhoon-2 project, which uses Trafilatura Barbaresi (2021) and text-dedup Mou et al. (2023) directly while creating its own heuristic filtering, to customizing the whole pipeline to specifically handle Thai. An example of a dedicated Thai data cleaning pipeline is from the OpenThaiGPT Yuenyong et al. (2025), in which they build the data preprocessing pipeline from scratch. However, these projects focus on open models rather than the openness of their training data and collection methods. While this provides a valuable resource, it means the pre-training corpora remain inaccessible and the pipeline’s design choices and empirical backing are not detailed, which is an important consideration for researchers focused on reproducibility. This raises a central question: What does it take to build a high-quality pre-training corpus for a language like Thai, and how should existing pipelines be adapted to meet its linguistic and cultural specifics?
@@ 4.3 Quality Filters
To justify the higher minimum document length threshold, we compared WangchanBERTa Lowphansirikul et al. (2021) perplexity statistics at lower bounds of 50 versus 171 words and found both mean and median perplexities to be lower at the higher threshold; therefore, we set the bound at 200 words. Finally, to speed up processing, we employ the ICU tokenizer rather than the slower nlpO3 tokenizer.
@@ 5.1 CPT and SFT Details
After CPT, we conduct supervised fine-tuning (SFT) using QLoRA Dettmers et al. (2023) on the Thai instruction dataset, named Wangchan-FLAN-6M [^11] https://huggingface.co/datasets/airesearch/WangchanX-FLAN-v6.1. We call the resultant model WangchanLION-V3-8B. For comparison, we employ Llama 3.1 8B base, SEA-LION-V3-8B-base, and Typhoon V2 8B base with the same SFT setting as our model. For consistency, all models have 8 billion parameters and use Llama-3.1-base as the original base model.
@@ 6.1.4 Using our data cleaning pipeline with cleaned data (FineWeb2)
Next, we question our data pipeline: can our data cleaning pipeline improve the data that has already been cleaned, like FineWeb2? To answer this question, we compare the original FineWeb2 and FineWeb2 processed with our pipeline. Our findings indicate that the original FineWeb2 performs slightly better overall than the version further cleaned with our pipeline in the Thai LLM Benchmark evaluation, as shown in Table 3. 8FineWeb2 processed with our pipeline substantially outperforms the original FineWeb2 on average in the SEA-HELM evaluation, as shown in Table 3. Moreover, we also found that, before applying our data pipeline, FineWeb2 had unwanted texts, e.g., low-quality, duplication of URLs and text overlap, adult content, gambling content, and PII. Further processed with our pipeline, we can significantly reduce its size from 35.9 million documents to 17.1 million, as shown in Table 4. This result is achieved by further cleaning and filtering data from FineWeb2. This implies that our benchmark can improve data quality by filtering low-quality data from the cleaned data. In addition, this also implies that the original cleaning data pipeline from FineWeb2 might not be applicable or suitable for Thai texts.
@@ 7 Conclusion and Outlook
Computational expense is still the main obstacle to studying the impact of data pipeline design. For this project, we secured 2,000 GPU-hours on H100 hardware, giving us effectively one full trial. Until reliable low-compute pre-training methods emerge, maximizing the public value of every costly run through complete knowledge sharing is essential to the collective progress of this field.
@@ Appendix B Language Identification
We conducted our experiments on a subsample of the raw Common Crawl dataset, consisting of the first 200,000 documents. Each of the three language identifiers was run separately on this sample to generate a language confidence score for each document. We used a threshold of 0.5—the same cutoff employed in the Dolma English corpus—to determine language inclusion. Based on this threshold, the number of documents retained by each language identifier was as follows:
@@
As illustrated in the Figure 4, fastText assigns high language confidence scores—often approaching 100%—to documents where the proportion of Thai characters, vowels, and intonation marks does not exceed 40%. This indicates a tendency of fastText to exhibit overconfidence in its language identification, potentially misclassifying documents with minimal Thai-specific character content. Notably, 8,008 documents (accounting for 4% of the 200,000-document sample) were identified by fastText as Thai with confidence scores above 0.5, despite containing 40% or fewer Thai-specific characters. Below is an example of such a document, demonstrating a high confidence score despite a low proportion of Thai characters.
To this end, it might seem that the only language identifier good enough for the Thai language is Lingua. However, considering Thai’s unique linguistic system and distinctive script, we concluded that relying on the Lingua language identifier may not be optimal. Therefore, we developed the ThaiCharRatioTagger, a custom language identifier that calculates the percentage of Thai characters, vowels, and intonation marks within a document. This approach simplifies the codebase and offers a more straightforward alternative to complex machine learning models.
@@ Appendix C Gopher Rules
In the original work, the authors noted that stacking quality filtering, content filtering, and deduplication results in a positive compounding effect; these rules overlap very little in the texts they remove. Our study comes to a similar conclusion regarding the Gopher rules. As shown in Figure 6 (Spearman correlation heatmap) and Figure 6 (Pearson correlation heatmap), we observe only low to moderate correlations between most rules. Some exceptions exist; for example, the "median word length" rule and the "fraction of words with alpha characters" rule are somewhat correlated. This is expected, as a higher median word length generally implies a higher number of alphabetic characters. A similar relationship is observed between the "word count" rule and the "required word count" rule. Based on this analysis, we choose to individually adjust the threshold for each rule. Note that for the rules "fraction of characters in most common n-gram" and "fraction of characters in duplicate n-grams," we use the average value within each group in our analysis, as similar rules tend to exhibit high intra-group correlation.

### v2

@@ 1 Introduction
For language-specific datasets, many data collection projects utilize existing data cleaning pipelines by customizing them to make these pipelines more suitable for their target languages. This ranges from making a small modification in one process, for instance, the Latxa project Etxaniz et al. (2024) for Basque, which adapted the Dolma and Corpus Cleaner v2 pipelines with a specific change in the filtering step, to customizing the whole pipeline. Examples of more extensive customization include the FinGPT Luukkonen et al. (2023) and IndicLLMSuite Khan et al. (2024) projects that built their own dedicated pipelines for Finnish and Indian languages, respectively. For the Thai language, the approach ranges from using a standard language-agnostic pipeline, such as the Typhoon-2 project, which uses Trafilatura Barbaresi (2021) and text-dedup Mou et al. (2023) directly while creating its own heuristic filtering, to customizing the whole pipeline to specifically handle Thai. An example of a dedicated Thai data cleaning pipeline is from the OpenThaiGPT Yuenyong et al. (2025), in which they build the data preprocessing pipeline from scratch. However, these projects focus on open models rather than the openness of their training data and collection methods. While this provides a valuable resource, it means the pre-training corpora remain inaccessible and the pipeline’s design choices and empirical backing are not detailed, which is an important consideration for researchers focused on reproducibility. This raises a central question: What does it take to build a high-quality pre-training corpus for a language like Thai, and how should existing pipelines be adapted to meet its linguistic and cultural specifics?
@@ 4.3 Quality Filters
To justify the higher minimum document length threshold, we compared WangchanBART 111111airesearch/wangchanbart-base perplexity statistics at lower bounds of 50 versus 171 words and found both mean and median perplexities to be lower at the higher threshold; therefore, we set the bound at 200 words. Finally, to speed up processing, we employ the ICU tokenizer rather than the slower nlpO3 tokenizer.
@@ 5.1 CPT and SFT Details
After CPT, we conduct supervised fine-tuning (SFT) using QLoRA Dettmers et al. (2023) on the Thai instruction dataset, named Wangchan-FLAN-6M [^12] https://huggingface.co/datasets/airesearch/WangchanX-FLAN-v6.1. We call the resultant model WangchanLION-V3-8B. For comparison, we employ Llama 3.1 8B base, SEA-LION-V3-8B-base, and Typhoon 2 8B base with the same SFT setting as our model. For consistency, all models have 8 billion parameters and use Llama-3.1-base as the original base model.
@@ 6.1.4 Using our data cleaning pipeline with cleaned data (FineWeb2)
Next, we question our data pipeline: can our data cleaning pipeline improve the data that has already been cleaned, like FineWeb2? To answer this question, we compare the original FineWeb2 and FineWeb2 processed with our pipeline. Our findings indicate that the original FineWeb2 performs slightly better overall than the version further cleaned with our pipeline in the Thai LLM Benchmark evaluation, as shown in Table 3. FineWeb2 processed with our pipeline substantially outperforms the original FineWeb2 on average in the SEA-HELM evaluation, as shown in Table 3. Moreover, we also found that, before applying our data pipeline, FineWeb2 had unwanted texts, e.g., low-quality, duplication of URLs and text overlap, adult content, gambling content, and PII. Further processed with our pipeline, we can significantly reduce its size from 35.9 million documents to 17.1 million, as shown in Table 4. This result is achieved by further cleaning and filtering data from FineWeb2. This implies that our benchmark can improve data quality by filtering low-quality data from the cleaned data. In addition, this also implies that the original cleaning data pipeline from FineWeb2 might not be applicable or suitable for Thai texts.
@@ 7 Conclusion and Outlook
Computational expense is still the main obstacle to studying the impact of data pipeline design. For this project, we secured 2,000 GPU-hours on H100 hardware, giving us effectively one full trial. Until reliable low-compute pre-training methods emerge, maximizing the public value of every costly run through complete knowledge sharing is essential to the collective progress of this field. With this limited experimentation budget, our model attains strong overall scores (approximately 2 points above recent Thai baselines), confirming the effectiveness of Thai-aware curation. The central contribution, however, is not the lead itself but the openly released pipeline, corpus, and checkpoints that let others build on and quickly surpass these results.
Acknowledgement
This research is supported by the National Research Foundation, Singapore, under its National Large Language Models Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore.
@@ Appendix B Language Identification
We conducted our experiments on a subsample of the raw Thai Common Crawl dataset, consisting of the first 200,000 documents. Each of the three language identifiers was run separately on this subsample to generate a language confidence score for each document. We used a threshold of 0.5, the same cut-off used in the Dolma English corpus, to determine language inclusion. Based on this threshold, the number of documents retained by each language identifier is shown in Table 7.
@@
As illustrated in the Figure 4, fastText assigns high language confidence scores—often approaching 100%—to documents where the proportion of Thai characters, vowels, and intonation marks does not exceed 40%. This indicates a tendency of fastText to exhibit overconfidence in its language identification, potentially misclassifying documents with minimal Thai-specific character content. Notably, 8,008 documents (accounting for 4% of the 200,000-document sample) were identified by fastText as Thai with confidence scores above 0.5, despite containing 40% or fewer Thai-specific characters.
To this end, it might seem that the only language identifier good enough for the Thai language is Lingua. However, considering Thai’s distinctive script, we concluded that relying on the Lingua language identifier may not be optimal. Therefore, we developed the ThaiCharRatioTagger, a custom language identifier that calculates the percentage of Thai characters, vowels, and intonation marks within a document. This approach simplifies the codebase and offers a more straightforward alternative to complex machine learning models.
@@ Appendix C Gopher Rules
In the Dolma work, the authors noted that stacking quality filtering, content filtering, and deduplication results in a positive compounding effect; these rules overlap very little in the texts they remove. Our study comes to a similar conclusion regarding the Gopher rules. As shown in Figure 6 (Spearman correlation heatmap) and Figure 6 (Pearson correlation heatmap), we observe only low to moderate correlations between most rules. Some exceptions exist; for example, the "median word length" rule and the "fraction of words with alpha characters" rule are somewhat correlated. This is expected, as a higher median word length generally implies a higher number of alphabetic characters. A similar relationship is observed between the "word count" rule and the "required word count" rule. Based on this analysis, we choose to individually adjust the threshold for each rule. Note that for the rules "fraction of characters in most common n-gram" and "fraction of characters in duplicate n-grams," we use the average value within each group in our analysis, as similar rules tend to exhibit high intra-group correlation.